OpenStudy (anonymous):

y=2x+1 y=-x+4 enter the answer in a ordered pair

pooja195 (pooja195):

E1-y=2x+1 E2-y=x+4 Use substituion they tell us what y equals right? SO sub E1 into E2 :) 2x+1=x+4 <---now solve for x

OpenStudy (anonymous):

the question is solve the equation by graphing enter your answer in a ordered pair i need help

pooja195 (pooja195):



pooja195 (pooja195):

See the intersection point? Thats the answer, sonow just make the coordinate